OpenStudy (anonymous):

with only % 7bills and $11bills avaliable it is impossible to make a stack of money worth exactley $3 it also impossible to make a stack worth $ 15 however it is possible to make a stack of money worth $ 25 by using two $ 7 and one $ 11 bill * Find all amounts that are impossible to make out of 7 and or 11 bills

hero (hero):

I believe you meant for your %7 to be a $7

OpenStudy (anonymous):

Yes do you get it

hero (hero):

11+ 7 = 18 will work 11+ 11 = 22 will work 7 + 7 = 14 will work 11*7 = 77 will work Multiples of 11 will work Multiples of seven will work Multiples of 18 will work Multiples of (11+7n) will work Multiples of (11n + 7) will work Multiples of (11n+7m) will work Other than those, any other combination will be "impossible".